# Sri Lanka Document Datasets: A Large-Scale, Multilingual Resource for Law, News, and Policy

### v3

@@ Abstract
We present a collection of open, machine-readable document datasets covering parliamentary proceedings, legal judgments, government publications, news, and tourism statistics from Sri Lanka. The collection currently comprises of 230,091 documents (57.7 GB) across 24 datasets in Sinhala, Tamil, and English. The datasets are updated daily and mirrored on GitHub and Hugging Face. These resources aim to support research in computational linguistics, legal analytics, socio-political studies, and multilingual natural language processing. We describe the data sources, collection pipeline, formats, and potential use cases, while discussing licensing and ethical considerations. This manuscript is at version v2025-10-16-0818.
@@ 3 Datasets
As of v2025-10-16-0818, Sri Lanka Document Datasets consists of 24 datasets and are publicly accessible on GitHub.  [^8] https://github.com/nuuuwan/lk_datasets
@@
    Hansard: A Hansard is the official verbatim record of parliamentary debates, preserving lawmakers’ words and decisions for history, law, and public accountability. 1,665 documents, 17.9 GB, from 2006-02-01 to 2025-09-24. Source: https://www.parliament.lk. Dataset: lk_hansard.
@@
    Appeal Court Judgements: A Court of Appeal judgment is a higher court ruling that reviews decisions of lower courts, shaping legal precedent and protecting citizens’ rights. 10,164 documents, 10.5 GB, from 2012-04-23 to 2025-10-15. Source: https://courtofappeal.lk. Dataset: lk_appeal_court_judgements.
@@
    Supreme Court Judgements: A Supreme Court judgment is a binding legal decision that interprets the Constitution and laws, shaping justice, governance, and citizens’ rights. 2,168 documents, 1.4 GB, from 2009-01-27 to 2025-10-15. Source: https://supremecourt.lk. Dataset: lk_supreme_court_judgements.
@@
    Police Press Releases: A police press release is an official update from law enforcement on crimes, arrests, safety alerts, or public notices, ensuring transparency and public awareness. 765 documents, 266.8 MB, from 2025-05-01 to 2025-10-15. Source: https://www.police.lk. Dataset: lk_police_press_releases.
@@
    Acts: A legal act is a law passed by Parliament that governs rights, duties, economy, and society, shaping daily life and national policy. 3,934 documents, 6.9 GB, from 1981-01-22 to 2025-10-07. Source: https://documents.gov.lk. Dataset: lk_acts.
@@
    Bills: A Bill is a draft law proposed in Parliament. It becomes binding once passed and enacted, shaping governance, rights, and daily life in the country. 4,080 documents, 1.9 GB, from 2010-05-10 to 2025-10-26. Source: https://documents.gov.lk. Dataset: lk_bills.
@@
    Extraordinary Gazettes 2020S: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 45,373 documents, 1.3 GB, from 2020-01-01 to 2025-10-15. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes_2020s.
@@
    Extraordinary Gazettes 2010S: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 56,379 documents, 3.3 GB, from 2010-01-01 to 2019-12-31. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes_2010s.
@@
    Cabinet Decisions: A Sri Lanka Cabinet Decision is an official policy or action agreed by the Cabinet of Ministers, shaping governance, law, and national development in the country. 10,385 documents, 136.4 MB, from 2010-09-27 to 2025-10-07. Source: https://www.cabinetoffice.gov.lk. Dataset: lk_cabinet_decisions.
@@
    Treasury Press Releases: A Sri Lanka Treasury press release shares key govt financial updates—on budgets, debt, or policy—vital for transparency, guiding investors, citizens, and markets on the nation’s economic direction. 134 documents, 144.5 MB, from 2015-09-08 to 2025-10-07. Source: https://www.treasury.gov.lk. Dataset: lk_treasury_press_releases.
@@
  12. 12.
    News: A collection of news documents. 81,155 documents, 1.2 GB, from 2021-09-12 to 2025-10-16. Source: multiple sources. Dataset: lk_news.
  13. 13.
    Tourism Weekly Reports: Report on Weekly Tourist Arrivals to Sri Lanka. 34 documents, 96.8 MB, from 2023-01-01 to 2025-10-01. Source: https://www.sltda.gov.lk. Dataset: lk_tourism_weekly_reports.
@@
    Tourism Monthly Reports: Report on Monthly Tourist Arrivals to Sri Lanka. 127 documents, 308.9 MB, from 2015-01-01 to 2025-08-01. Source: multiple sources. Dataset: lk_tourism_monthly_reports.
@@
    Dmc Situation Reports: Situation Report including information about Heavy Rain, Wind, Tree Falling, Lighting etc. 4,321 documents, 3.0 GB, from 2018-01-02 to 2025-10-15. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_situation_reports.
@@
    Dmc Weather Forecasts: Weather Forecasts for various places in Sri Lanka. 3,660 documents, 4.4 GB, from 2023-03-26 to 2025-10-16. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_weather_forecasts.
@@
    Dmc River Water Level And Flood Warnings: River Water Level and Flood Warnings for various places in Sri Lanka. 21 documents, 7.5 MB, from 2025-06-10 to 2025-10-15. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_river_water_level_and_flood_warnings.
@@
    Dmc Landslide Warnings: Landslide Warnings including early warnings, locations of potential risk, areas and places which need special attention, and automated landslide early warning map. 564 documents, 437.8 MB, from 2019-09-26 to 2025-10-13. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_landslide_warnings.
@@
  20. 20.
    Fisheries Annual Statistics Reports: Annual Fisheries Statistics Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 9 documents, 17.4 MB, from 2017-01-01 to 2024-01-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_annual_statistics_reports.
  21. 21.
    Fisheries Monthly Fish Production Reports: Monthly Fish Production Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 96 documents, 10.3 MB, from 2019-01-01 to 2025-08-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_monthly_fish_production_reports.
@@
    Fisheries Weekly Fish Prices Reports: Weekly Fish Prices Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 248 documents, 22.1 MB, from 2019-01-01 to 2025-09-22. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_weekly_fish_prices_reports.
@@
    Fisheries Monthly Export Import Reports: Monthly Fish Export and Import Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 64 documents, 51.6 MB, from 2019-01-01 to 2025-08-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_monthly_export_import_reports.
@@
    Edupub: Educational Publications from the Educational Publications Department, Sri Lanka. 1,426 documents, 866.2 MB, from 2025-01-01 to 2025-01-01. Source: http://www.edupub.gov.lk. Dataset: lk_edupub.

### v4

@@ Abstract
We present a collection of open, machine-readable document datasets covering parliamentary proceedings, legal judgments, government publications, news, and tourism statistics from Sri Lanka. The collection currently comprises of 247,818 documents (67.6 GB) across 26 datasets in Sinhala, Tamil, and English. The datasets are updated daily and mirrored on GitHub and Hugging Face. These resources aim to support research in computational linguistics, legal analytics, socio-political studies, and multilingual natural language processing. We describe the data sources, collection pipeline, formats, and potential use cases, while discussing licensing and ethical considerations. This manuscript is at version v2026-01-03-0933.
@@ 3 Datasets
As of v2026-01-03-0933, Sri Lanka Document Datasets consists of 26 datasets and are publicly accessible on GitHub.  [^8] https://github.com/nuuuwan/lk_datasets
@@
    Hansard 2020S: A Hansard is the official verbatim record of parliamentary debates, preserving lawmakers’ words and decisions for history, law, and public accountability. 188 documents, 2.5 GB, from 2023-11-17 to 2025-11-25. Source: https://www.parliament.lk. Dataset: lk_hansard_2020s.
@@
    Hansard 2010S: A Hansard is the official verbatim record of parliamentary debates, preserving lawmakers’ words and decisions for history, law, and public accountability. 10 documents, 153.2 MB, from 2019-09-05 to 2019-11-11. Source: https://www.parliament.lk. Dataset: lk_hansard_2010s.
@@
    Hansard 2000S: A Hansard is the official verbatim record of parliamentary debates, preserving lawmakers’ words and decisions for history, law, and public accountability. 10 documents, 135.4 MB, from 2009-10-20 to 2009-12-08. Source: https://www.parliament.lk. Dataset: lk_hansard_2000s.
@@
    Appeal Court Judgements: A Court of Appeal judgment is a higher court ruling that reviews decisions of lower courts, shaping legal precedent and protecting citizens’ rights. 10,351 documents, 10.5 GB, from 2012-04-23 to 2025-12-NA. Source: https://courtofappeal.lk. Dataset: lk_appeal_court_judgements.
@@
    Supreme Court Judgements: A Supreme Court judgment is a binding legal decision that interprets the Constitution and laws, shaping justice, governance, and citizens’ rights. 2,449 documents, 1.5 GB, from 2009-01-27 to 2025-12-19. Source: https://supremecourt.lk. Dataset: lk_supreme_court_judgements.
@@
    Police Press Releases: A police press release is an official update from law enforcement on crimes, arrests, safety alerts, or public notices, ensuring transparency and public awareness. 1,021 documents, 345.4 MB, from 2025-05-01 to 2025-12-30. Source: https://www.police.lk. Dataset: lk_police_press_releases.
@@
    Acts: A legal act is a law passed by Parliament that governs rights, duties, economy, and society, shaping daily life and national policy. 3,949 documents, 6.8 GB, from 1981-01-22 to 2025-12-26. Source: https://documents.gov.lk. Dataset: lk_acts.
@@
    Bills: A Bill is a draft law proposed in Parliament. It becomes binding once passed and enacted, shaping governance, rights, and daily life in the country. 4,140 documents, 1.8 GB, from 2010-05-10 to 2025-12-29. Source: https://documents.gov.lk. Dataset: lk_bills.
@@
    Extraordinary Gazettes 2020S: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 47,123 documents, 9.3 GB, from 2020-01-01 to 2025-12-31. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes_2020s.
@@
    Extraordinary Gazettes 2010S: An Extraordinary Gazette is an official government publication used to announce urgent laws, regulations, or public notices with immediate effect. 56,379 documents, 14.4 GB, from 2010-01-01 to 2019-12-31. Source: https://documents.gov.lk. Dataset: lk_extraordinary_gazettes_2010s.
@@
    Cabinet Decisions: A Sri Lanka Cabinet Decision is an official policy or action agreed by the Cabinet of Ministers, shaping governance, law, and national development in the country. 10,569 documents, 137.7 MB, from 2010-09-27 to 2025-12-29. Source: https://www.cabinetoffice.gov.lk. Dataset: lk_cabinet_decisions.
  12. 12.
    Treasury Press Releases: A Sri Lanka Treasury press release shares key govt financial updates—on budgets, debt, or policy—vital for transparency, guiding investors, citizens, and markets on the nation’s economic direction. 140 documents, 150.9 MB, from 2015-09-08 to 2025-12-19. Source: https://www.treasury.gov.lk. Dataset: lk_treasury_press_releases.
  13. 13.
@@
    News: A collection of news documents. 96,709 documents, 1.5 GB, from 2021-09-12 to 2026-01-03. Source: multiple sources. Dataset: lk_news.
@@
    Tourism Weekly Reports: Report on Weekly Tourist Arrivals to Sri Lanka. 36 documents, 106.1 MB, from 2023-01-01 to 2025-12-01. Source: https://www.sltda.gov.lk. Dataset: lk_tourism_weekly_reports.
@@
    Tourism Monthly Reports: Report on Monthly Tourist Arrivals to Sri Lanka. 130 documents, 335.4 MB, from 2015-01-01 to 2025-11-01. Source: multiple sources. Dataset: lk_tourism_monthly_reports.
@@
    Dmc Situation Reports: Situation Report including information about Heavy Rain, Wind, Tree Falling, Lighting etc. 4,425 documents, 3.0 GB, from 2018-01-02 to 2026-01-02. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_situation_reports.
@@
    Dmc Weather Forecasts: Weather Forecasts for various places in Sri Lanka. 4,055 documents, 5.0 GB, from 2023-03-26 to 2026-01-03. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_weather_forecasts.
@@
    Dmc River Water Level And Flood Warnings: River Water Level and Flood Warnings for various places in Sri Lanka. 326 documents, 128.9 MB, from 2025-06-10 to 2026-01-01. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_river_water_level_and_flood_warnings.
  20. 20.
    Dmc Landslide Warnings: Landslide Warnings including early warnings, locations of potential risk, areas and places which need special attention, and automated landslide early warning map. 630 documents, 479.1 MB, from 2019-09-26 to 2025-12-23. Source: https://www.dmc.gov.lk. Dataset: lk_dmc_landslide_warnings.
  21. 21.
@@
    Fisheries Annual Statistics Reports: Annual Fisheries Statistics Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 9 documents, 15.1 MB, from 2017-01-01 to 2024-01-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_annual_statistics_reports.
@@
    Fisheries Monthly Fish Production Reports: Monthly Fish Production Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 98 documents, 9.2 MB, from 2019-01-01 to 2025-10-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_monthly_fish_production_reports.
@@
    Fisheries Weekly Fish Prices Reports: Weekly Fish Prices Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 261 documents, 22.9 MB, from 2019-01-01 to 2025-12-22. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_weekly_fish_prices_reports.
  25. 25.
    Fisheries Monthly Export Import Reports: Monthly Fish Export and Import Reports of the Ministry of Fisheries,Aquatic and Ocean Resources, Sri Lanka 65 documents, 51.4 MB, from 2019-01-01 to 2025-09-01. Source: https://www.fisheries.gov.lk. Dataset: lk_fisheries_monthly_export_import_reports.
  26. 26.
    Edupub: Educational Publications from the Educational Publications Department, Sri Lanka. 1,426 documents, 5.6 GB, from 2025-01-01 to 2025-01-01. Source: http://www.edupub.gov.lk. Dataset: lk_edupub.